# Study of semileptonic $B\to DP\ell^{+}\nu_{\ell}$ decays based on the SU(3) flavor symmetry

### v1

@@ III.2 Non-leptonic two-body $D_{J}\to DP$ decays
Non-leptonic two-body $D_{J}\to DP$ decays via strong or electromagnetic interaction are presented in this section. Due to parity conservation, only $D_{J}=D_{0}/D^{*}/D^{*}_{2}$ resonances are considered for the $B\to DP\ell^{+}\nu_{\ell}$ decays Kim:1998np . In terms of the SU(3) flavor symmetry, the decay amplitudes of the strong or electromagnetic $D_{J}\to DP$ decays can be parameterized as
@@
where $a^{D_{J}}_{01,02}$ are the nonperturbative coefficients, and all $D_{J}$ are two-quark state. $a^{D_{J}}_{02}$ are OZI suppressed and it will be ignored in later numerical analysis. The decay amplitudes for each $D^{*}_{2}\to DP$ decay are summarized in Tab. 7. The decay amplitudes for $D^{*}/D_{0}\to DP$ can be obtained by replacing $a^{D^{*}_{2}}_{01,02}$ listed in Tab. 7 with $a^{D^{*},D_{0}}_{01,02}$ only if their phase spaces are allowed.
@@
For the $D^{*}_{2}\to DP$ decays, since there is no experimental data of the branching ratios, we can not constrain $a^{D^{*}_{2}}_{01}$ directly. Nevertheless, $\frac{\mathcal{B}(D^{*}_{2}\to D\pi)}{\mathcal{B}(D^{*}_{2}\to D^{*}\pi)}=1.52\pm 0.14$ within $2\sigma$ errors are measured ParticleDataGroup:2022pth . The SU(3) relation of the decay amplitudes of the $D^{*}_{2}\to D^{*}P$ decays are given in Ref. Wan:2024ssl . Using $\frac{\mathcal{B}(D^{*}_{2}\to D\pi)}{\mathcal{B}(D^{*}_{2}\to D^{*}\pi)}=1.52\pm 0.14$ and assuming $\mathcal{B}(D^{*0}_{2}\to D\pi)+\mathcal{B}(D^{*0}_{2}\to D^{*}\pi)\leq 1$, $\mathcal{B}(D^{*+}_{2}\to D\pi)+\mathcal{B}(D^{*+}_{2}\to D^{*}\pi)\leq 1$ and $\mathcal{B}(D^{*+}_{s2}\to DK)+\mathcal{B}(D^{*+}_{s2}\to D^{*}K)\leq 1$, one can constrain the nonperturbative coefficients $a^{D^{*}_{2}}_{01}$, and we obtain that $|a^{D^{*}_{2}}_{01}|=25.14\pm 1.47$. Then the branching ratios of the $D^{*}_{2}\to DP$ decays can be predicted, which are summarized in the second column of Tab. 7. In addition, their decay width predictions and previous width predictions are also given in the forth and fifth columns of Tab. 7, respectively. Our width predictions are about 1 time larger than ones in Ref. Wang:2014oca .
@@
$D^{*0}_{2}(2460)\to D^{0}\pi^{0}$ $a^{D^{*}_{2}}_{01}/\sqrt{2}$ $D^{*+}_{2}(2460)\to D^{+}\pi^{0}$ $-a^{D^{*}_{2}}_{01}/\sqrt{2}$
$D^{*0}_{2}(2460)\to D^{0}\eta$ $a^{D^{*}_{2}}_{01}\Big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\Big{)}-\sqrt{3}a^{D^{*}_{2}}_{02}sin\theta_{P}$ $D^{*+}_{2}(2460)\to D^{+}\eta$ $a^{D^{*}_{2}}_{01}\Big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\Big{)}-\sqrt{3}a^{D^{*}_{2}}_{02}sin\theta_{P}$
$D^{*0}_{2}(2460)\to D^{+}\pi^{-}$ $a^{D^{*}_{2}}_{01}$ $D^{*+}_{2}(2460)\to D^{0}\pi^{+}$ $a^{D^{*}_{2}}_{01}$
$D^{*0}_{2}(2460)\to D^{+}_{s}K^{-}$ $a^{D^{*}_{2}}_{01}$ $D^{*+}_{2}(2460)\to D^{+}_{s}\overline{K}^{0}$ $a^{D^{*}_{2}}_{01}$
$D^{*+}_{s2}(2573)\to D^{0}K^{+}$ $a^{D^{*}_{2}}_{01}$
$D^{*+}_{s2}(2573)\to D^{+}K^{0}$ $a^{D^{*}_{2}}_{01}$
$D^{*+}_{s2}(2573)\to D^{+}_{s}\eta$ $-a^{D^{*}_{2}}_{01}\Big{(}\frac{2cos\theta_{P}}{\sqrt{6}}+\frac{sin\theta_{P}}{\sqrt{3}}\Big{)}-\sqrt{3}a^{D^{*}_{2}}_{02}sin\theta_{P}$
Table 7: The predictions of the $D^{*}_{2}\to DP$ decays within $2\sigma$ errors.
Decay modes Branching ratios (×$(\times 10^{-2})$ Decay widthes (MeV) Decay widthes from Wang:2014oca (MeV)
$D^{*0}_{2}(2460)\to D^{0}\pi^{0}$ $20.18\pm 1.77$ $9.59\pm 1.06$ $4.14^{+1.82}_{-1.57}$
$D^{*0}_{2}(2460)\to D^{0}\eta$ $0.13\pm 0.03$ $0.06\pm 0.02$cdots
$D^{*0}_{2}(2460)\to D^{+}\pi^{-}$ $38.51\pm 3.41$ $18.30\pm 2.03$ $7.91^{+3.49}_{-3.00}$
$D^{*0}_{2}(2460)\to D^{+}_{s}K^{-}$ $[1.82\times 10^{-15},6.56\times 10^{-6}]$ $[8.77\times 10^{-14},3.00\times 10^{-6}]$cdots
$D^{*+}_{2}(2460)\to D^{+}\pi^{0}$ $19.40\pm 1.63$ $9.33\pm 0.92$cdots
$D^{*+}_{2}(2460)\to D^{+}\eta$ $0.11\pm 0.04$ $(5.45\pm 2.15)\times 10^{-2}$cdots
$D^{*+}_{2}(2460)\to D^{0}\pi^{+}$ $39.88\pm 3.33$ $19.18\pm 1.90$cdots
$D^{*+}_{2}(2460)\to D^{+}_{s}\overline{K}^{0}$ $[1.26\times 10^{-14},1.80\times 10^{-6}]$ $[6.01\times 10^{-13},8.28\times 10^{-5}]$cdots
$D^{*+}_{s2}(2573)\to D^{0}K^{+}$ $42.32\pm 5.20$ $7.52\pm 0.95$ $3.35^{+1.48}_{-1.27}$
$D^{*+}_{s2}(2573)\to D^{+}K^{0}$ $38.12\pm 4.72$ $6.79\pm 0.86$ $3.04^{+1.34}_{-1.15}$
$D^{*+}_{s2}(2573)\to D^{+}_{s}\eta$ $0.58\pm 0.19$ $(10.23\pm 3.43)\times 10^{-2}$cdots
@@ III.3 Numerical results of the resonant $B\to DP\ell^{+}\nu_{\ell}$ decays
In terms of $\mathcal{B}(B\rightarrow D_{J}\ell^{+}\nu_{\ell})$ given in Tabs. 5-5 and $\mathcal{B}(D_{J}\rightarrow DP)$ given in Eqs. (52-53) and Tab. 7, after considering the further experimental bounds of the resonant $B\to D\pi\ell^{+}\nu_{\ell}$ decays given in Eqs. (2-3) and (5-6), one can obtain the branching ratio predictions of the resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, and they are listed in the third, forth and fifth columns of Tab. 2 for the $D^{*}$, $D_{0}$ and $D_{(s)2}^{*}$ resonances, respectively. Corresponding experimental data with $2\sigma$ errors are also listed in Tab. 2 for the convenience of comparison. Note that, since the vector resonances are also considered in this work, $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ and $\mathcal{B}(B^{0}\to D^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ in Eq. (1) and Eq. (4) are not used for our results. Many resonant branching ratios are predicted for the first time.
@@
All current experimental data of $\mathcal{B}(B\to DP\ell^{+}\nu_{\ell})$ can be explained by the SU(3) flavor symmetry approach. For the SU(3) flavor breaking effects in the $B\to DP\ell^{+}\nu_{\ell}$ decays, which are in the similar situation to the $B\to D^{*}P\ell^{+}\nu_{\ell}$ decays. The SU(3) flavor breaking effects due to a nonzero $s$ quark mass dominate the SU(3) breaking effects. As given in Tab. 2, the dominate SU(3) flavor breaking effects might appear in the non-resonant and the charmed tensor resonant $B\to DP\ell^{+}\nu_{\ell}$ decays. Nevertheless, there is only the data for $\mathcal{B}(B_{u,d}\to D\pi\ell^{+}\nu_{\ell})_{D^{*}_{2}}$, and there is not any data for $\mathcal{B}(B_{s}\to DK\ell^{+}\nu_{\ell})_{D^{*}_{s2}}$. Or there is only the data of $\mathcal{B}(B^{+}\to D_{s}^{-}K^{+}\ell^{+}\nu_{\ell})_{N}$, and there is no any data of $\mathcal{B}(B_{u,d,s}\to DK\ell^{+}\nu_{\ell})_{N}$. Therefore, we can not directly judge how large the possible SU(3) breaking effects are in the $B\to DP\ell^{+}\nu_{\ell}$ decays.
@@
where $d\mathcal{B}(B\to D^{*}(\lambda)\ell^{+}\nu_{\ell},t_{V})/dq^{2}$ and $\mathcal{B}(D^{*}(\lambda)\to D\pi,t_{V})$ are obtained from to Eq. (44) and Eq. (50) by replacing $m_{D^{*}}\to\sqrt{t_{V}}$, respectively. There are two nonperturbative coefficients $V^{BD^{*}}(0)$ in $d\mathcal{B}(B\to D^{*}(\lambda)\ell^{+}\nu_{\ell},t_{V})/dq^{2}$ and $a_{01}^{D^{*}}$ in $\mathcal{B}(D^{*}(\lambda)\to D\pi,t_{V})$. $V^{BD^{*}}(0)=0.65\pm 0.05$ from the data of $\mathcal{B}(B\to D^{*}\ell^{+}\nu_{\ell})$ listed in Tab. 5, and $a_{01}^{D^{*}}=8.42\pm 0.38$ from the data of $\mathcal{B}(D^{*+}\to D\pi)$ given in Eq. (52). $\Gamma_{D^{*0}}=(56.00\pm 5.74)\leavevmode\nobreak\ KeV$, which is obtained via the experimental data of $\mathcal{B}(D^{*0}\to D^{0}\pi^{0})$ in Eq. (52) and the SU(3) flavor symmetry in $D^{*}\to D\pi$. Following Refs. Cheng:1993ah , choosing $n=3$, the results of $\mathcal{B}(B\to D\pi\ell^{+}\nu_{\ell})_{D^{*}}$ are obtained, and they are listed in Tab. by denoting $\dagger$. One can see that the results obtained by considering the $D^{*}$ width effects are slightly smaller than ones obtained by the narrow width approximation.
@@ IV Summary
The semileptonic $B\to DP\ell^{+}\nu_{\ell}$ decays with the non-resonances, the vector resonances, the scalar resonances and the tensor resonances have been investigated by using the SU(3) flavor symmetry based on the relevant experimental data. The amplitude relations of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, the semileptonic $B\to D/D^{*}/D_{0}/D^{*}_{2}\ell^{+}\nu_{\ell}$ decays and the non-leptonic $D^{*}/D_{0}/D^{*}_{2}\to DP$ decays have been obtained, and then the resonant branching ratios have been obtained by the narrow width approximation after considering the resonant experimental data of the $B\to DP\ell^{\prime+}\nu_{\ell^{\prime}}$ decays. Our main results can be summarized as follows.
For the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, the central values of $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}}$ $\mathcal{B}(B^{0}\to D^{-}_{s}K^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}_{s}\to\overline{D}^{0}K^{-}\ell^{\prime+}\nu_{\ell^{\prime}}$ $\mathcal{B}(B^{0}_{s}\to D^{-}\overline{K}^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ and $\mathcal{B}(B^{0}_{s}\to D_{s}^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ are on the orders of $10^{-4}$, which could be measured by the LHCb and Belle II experiments. Other non-resonant decay branching ratios of $B^{+}\to\overline{D}^{0}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, $B^{0}\to D^{-}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$, $B^{0}_{s}\to D_{s}^{-}\eta^{\prime}\ell^{\prime+}\nu_{\ell^{\prime}}$ and all $B\to DP\tau^{+}\nu_{\tau}$ decays are strongly suppressed by the narrow phase spaces.
@@
Although only approximate predictions can be obtained by the SU(3) flavor symmetry, they are still useful for understanding these decays. Until now, our predictions of the $B\to DP\ell^{+}\nu_{\ell}$ decays are quite coincident with present experimental data, and they could be tested in future experiments, such as LHCb and Belle II.

### v2

@@ III.2 Decays $D_{J}\to DP$
Non-leptonic two-body $D_{J}\to DP$ decays via strong or electromagnetic interaction are presented in this section. In terms of the SU(3) flavor symmetry, the decay amplitudes of the strong or electromagnetic $D_{J}\to DP$ decays can be parameterized as
@@
where $a^{D_{J}}_{01,02}$ are the non-perturbative coefficients, and all $D_{J}$ are two-quark states. $a^{D_{J}}_{02}$ are OZI suppressed and it will be ignored in later numerical analysis. The decay amplitudes for each $D^{*}_{2}\to DP$ decay are summarized in Tab. 7. The decay amplitudes for $D^{*}/D_{0}\to DP$ can be obtained by replacing $a^{D^{*}_{2}}_{01,02}$ listed in Tab. 7 with $a^{D^{*},D_{0}}_{01,02}$ only if their phase spaces are allowed.
@@
For the $D^{*}_{2}\to DP$ decays, since there is no experimental data of the branching ratios, we can not constrain $a^{D^{*}_{2}}_{01}$ directly. Nevertheless, $\frac{\mathcal{B}(D^{*}_{2}\to D\pi)}{\mathcal{B}(D^{*}_{2}\to D^{*}\pi)}=1.52\pm 0.14$ within $2\sigma$ errors are measured ParticleDataGroup:2022pth . The SU(3) relation of the decay amplitudes of the $D^{*}_{2}\to D^{*}P$ decays are given in Ref. Wan:2024ssl . Using $\frac{\mathcal{B}(D^{*}_{2}\to D\pi)}{\mathcal{B}(D^{*}_{2}\to D^{*}\pi)}=1.52\pm 0.14$ and assuming $\mathcal{B}(D^{*0}_{2}\to D\pi)+\mathcal{B}(D^{*0}_{2}\to D^{*}\pi)\leq 1$, $\mathcal{B}(D^{*+}_{2}\to D\pi)+\mathcal{B}(D^{*+}_{2}\to D^{*}\pi)\leq 1$ and $\mathcal{B}(D^{*+}_{s2}\to DK)+\mathcal{B}(D^{*+}_{s2}\to D^{*}K)\leq 1$, one can constrain the non-perturbative coefficients $a^{D^{*}_{2}}_{01}$, and we obtain that $|a^{D^{*}_{2}}_{01}|=25.14\pm 1.47$. Then the branching ratios of the $D^{*}_{2}\to DP$ decays can be predicted, which are given in the second column of Tab. 7. In addition, their decay width predictions and previous width predictions are also given in the third and fourth columns of Tab. 7, respectively. Our width predictions are about 1 time larger than ones in Ref. Wang:2014oca , nevertheless, they are very close to ones in Ref. Zhang:2016dom .
@@
$D^{*0}_{2}\to D^{0}\pi^{0}$ $a^{D^{*}_{2}}_{01}/\sqrt{2}$ $D^{*+}_{2}\to D^{+}\pi^{0}$ $-a^{D^{*}_{2}}_{01}/\sqrt{2}$
$D^{*0}_{2}\to D^{0}\eta$ $a^{D^{*}_{2}}_{01}\Big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\Big{)}-\sqrt{3}a^{D^{*}_{2}}_{02}sin\theta_{P}$ $D^{*+}_{2}\to D^{+}\eta$ $a^{D^{*}_{2}}_{01}\Big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\Big{)}-\sqrt{3}a^{D^{*}_{2}}_{02}sin\theta_{P}$
$D^{*0}_{2}\to D^{+}\pi^{-}$ $a^{D^{*}_{2}}_{01}$ $D^{*+}_{2}\to D^{0}\pi^{+}$ $a^{D^{*}_{2}}_{01}$
$D^{*0}_{2}\to D^{+}_{s}K^{-}$ $a^{D^{*}_{2}}_{01}$ $D^{*+}_{2}\to D^{+}_{s}\overline{K}^{0}$ $a^{D^{*}_{2}}_{01}$
$D^{*+}_{s2}\to D^{0}K^{+}$ $a^{D^{*}_{2}}_{01}$
$D^{*+}_{s2}\to D^{+}K^{0}$ $a^{D^{*}_{2}}_{01}$
$D^{*+}_{s2}\to D^{+}_{s}\eta$ $-a^{D^{*}_{2}}_{01}\Big{(}\frac{2cos\theta_{P}}{\sqrt{6}}+\frac{sin\theta_{P}}{\sqrt{3}}\Big{)}-\sqrt{3}a^{D^{*}_{2}}_{02}sin\theta_{P}$
Table 7: The branching ratio predictions of the $D^{*}_{2}\to DP$ decays within $2\sigma$ errors.
Decay modes Branching ratios (×$(\times 10^{-2})$ Decay widthes (MeV) Decay widthes from (MeV)
$D^{*0}_{2}\to D^{0}\pi^{0}$ $20.18\pm 1.77$ $9.59\pm 1.06$ $4.14^{+1.82}_{-1.57}$ Wang:2014oca , $12.0$ Zhang:2016dom
$D^{*0}_{2}\to D^{0}\eta$ $0.13\pm 0.03$ $0.06\pm 0.02$cdots
$D^{*0}_{2}\to D^{+}\pi^{-}$ $38.51\pm 3.41$ $18.30\pm 2.03$ $7.91^{+3.49}_{-3.00}$ Wang:2014oca , $22.8$ Zhang:2016dom
$D^{*0}_{2}\to D^{+}_{s}K^{-}$ $[1.82\times 10^{-15},6.56\times 10^{-6}]$ $[8.77\times 10^{-14},3.00\times 10^{-6}]$cdots
$D^{*+}_{2}\to D^{+}\pi^{0}$ $19.40\pm 1.63$ $9.33\pm 0.92$cdots
$D^{*+}_{2}\to D^{+}\eta$ $0.11\pm 0.04$ $(5.45\pm 2.15)\times 10^{-2}$cdots
$D^{*+}_{2}\to D^{0}\pi^{+}$ $39.88\pm 3.33$ $19.18\pm 1.90$cdots
$D^{*+}_{2}\to D^{+}_{s}\overline{K}^{0}$ $[1.26\times 10^{-14},1.80\times 10^{-6}]$ $[6.01\times 10^{-13},8.28\times 10^{-5}]$cdots
$D^{*+}_{s2}\to D^{0}K^{+}$ $42.32\pm 5.20$ $7.52\pm 0.95$ $3.35^{+1.48}_{-1.27}$ Wang:2014oca , $9.49$ Zhang:2016dom
$D^{*+}_{s2}\to D^{+}K^{0}$ $38.12\pm 4.72$ $6.79\pm 0.86$ $3.04^{+1.34}_{-1.15}$ Wang:2014oca , $8.61$ Zhang:2016dom
$D^{*+}_{s2}\to D^{+}_{s}\eta$ $0.58\pm 0.19$ $(10.23\pm 3.43)\times 10^{-2}$cdots
@@ III.3 Numerical results of the resonant $B\to DP\ell^{+}\nu_{\ell}$ decays
In terms of $\mathcal{B}(B\rightarrow D_{J}\ell^{+}\nu_{\ell})$ given in Tabs. 5-5 and $\mathcal{B}(D_{J}\rightarrow DP)$ given in Eqs. (49-50) and Tab. 7, after considering the further experimental bounds of the resonant $B\to D\pi\ell^{+}\nu_{\ell}$ decays given in Eqs. (2-3) and (5-6), one can obtain the branching ratio predictions of the resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, and they are listed in the third, fourth and fifth columns of Tab. 2 for the $D^{*}$, $D_{0}$ and $D_{(s)2}^{*}$ resonances, respectively. Corresponding experimental data with $2\sigma$ errors are also listed in Tab. 2 for the convenience of comparison. Note that, since the vector resonances are also considered in this work, $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ and $\mathcal{B}(B^{0}\to D^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{T}$ in Eq. (1) and Eq. (4), which only conclude $D_{0}$ and $D^{*}_{2}$ resonances, are not used for our results. Many resonant branching ratios in Tab. 2 are predicted for the first time.
@@
All present experimental data of $\mathcal{B}(B\to DP\ell^{+}\nu_{\ell})$ may be explained by the SU(3) flavor symmetry approach. The SU(3) flavor breaking effects in the $B\to DP\ell^{+}\nu_{\ell}$ decays are in the similar to ones in the $B\to D^{*}P\ell^{+}\nu_{\ell}$ decays. As given in Tab. 2, the dominate SU(3) flavor breaking effects might appear in the non-resonant and the charmed tensor resonant $B\to DP\ell^{+}\nu_{\ell}$ decays. Nevertheless, there is only the data for $\mathcal{B}(B_{u,d}\to D\pi\ell^{+}\nu_{\ell})_{D^{*}_{2}}$, and there is not any data for $\mathcal{B}(B_{s}\to DK\ell^{+}\nu_{\ell})_{D^{*}_{s2}}$. Or there is only the data of $\mathcal{B}(B^{+}\to D_{s}^{-}K^{+}\ell^{+}\nu_{\ell})_{N}$, and there is no data of $\mathcal{B}(B_{u,d,s}\to DK\ell^{+}\nu_{\ell})_{N}$. Therefore, we can not directly judge how large the possible SU(3) breaking effects are in the $B\to DP\ell^{+}\nu_{\ell}$ decays.
@@
where $d\mathcal{B}(B\to D^{*}(\lambda)\ell^{+}\nu_{\ell},t_{V})/dq^{2}$ and $\mathcal{B}(D^{*}(\lambda)\to D\pi,t_{V})$ are obtained from to Eq. (28) and Eq. (47) by replacing $m_{D^{*}}\to\sqrt{t_{V}}$, respectively. There are two non-perturbative coefficients $V^{BD^{*}}(0)$ in $d\mathcal{B}(B\to D^{*}(\lambda)\ell^{+}\nu_{\ell},t_{V})/dq^{2}$ and $a_{01}^{D^{*}}$ in $\mathcal{B}(D^{*}(\lambda)\to D\pi,t_{V})$. $V^{BD^{*}}(0)=0.65\pm 0.05$ from the data of $\mathcal{B}(B\to D^{*}\ell^{+}\nu_{\ell})$ listed in Tab. 5, and $a_{01}^{D^{*}}=8.42\pm 0.38$ from the data of $\mathcal{B}(D^{*+}\to D\pi)$ given in Eq. (49). $\Gamma_{D^{*0}}=(56.00\pm 5.74)\leavevmode\nobreak\ KeV$, which is obtained via the experimental data of $\mathcal{B}(D^{*0}\to D^{0}\pi^{0})$ in Eq. (49) and the SU(3) flavor symmetry in $D^{*}\to D\pi$. Following Refs. Cheng:1993ah , choosing $n=3$, the results of $\mathcal{B}(B\to D\pi\ell^{+}\nu_{\ell})_{D^{*}}$ are obtained, and they are listed in Tab. 2 by denoting $\dagger$. One can see that the results obtained by considering the $D^{*}$ width effects are slightly smaller than ones obtained by the narrow width approximation.
@@ IV Summary
The semileptonic $B\to DP\ell^{+}\nu_{\ell}$ decays with the non-resonances, the vector resonances, the scalar resonances and the tensor resonances have been explored in terms of the SU(3) flavor symmetry based on the relevant experimental data. The amplitude relations of the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, the semileptonic $B\to D/D^{*}/D_{0}/D^{*}_{2}\ell^{+}\nu_{\ell}$ decays and the non-leptonic $D^{*}/D_{0}/D^{*}_{2}\to DP$ decays have been obtained, and then the resonant branching ratios have been obtained by the narrow width approximation after considering the resonant experimental data of the $B\to DP\ell^{\prime+}\nu_{\ell^{\prime}}$ decays. Our main results can be summarized as follows.
For the non-resonant $B\to DP\ell^{+}\nu_{\ell}$ decays, the central values of $\mathcal{B}(B^{+}\to D^{-}\pi^{+}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{+}\to\overline{D}^{0}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to\overline{D}^{0}\pi^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ $\mathcal{B}(B^{0}\to D^{-}_{s}K^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\pi^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}\to D^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$, $\mathcal{B}(B^{0}_{s}\to\overline{D}^{0}K^{-}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ $\mathcal{B}(B^{0}_{s}\to D^{-}\overline{K}^{0}\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ and $\mathcal{B}(B^{0}_{s}\to D_{s}^{-}\eta\ell^{\prime+}\nu_{\ell^{\prime}})_{N}$ are on the orders of $10^{-4}$, which could be measured by the LHCb and Belle II experiments. Other non-resonant decays are strongly suppressed by the narrow phase spaces.
@@
Although the SU(3) flavor symmetry only can give the approximate predictions, they are still very useful for understanding these decays. Until now, our predictions of the $B\to DP\ell^{+}\nu_{\ell}$ decays are quite coincident with present experimental data, and our predictions could be tested in future experiments, such as LHCb and Belle II.